OpenStudy (anonymous):

Solving Radical equations square root 6x-8= x

OpenStudy (anonymous):

6x-x = 8

OpenStudy (anonymous):

How do you know it?

OpenStudy (anonymous):

What is the process to get the answer ?

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

just put like terms together to find their values

OpenStudy (anonymous):

square root 6x-8=x 6x-6x-8=x x-8=x x-x-8=x-x x=8 is that ? @sangya21

OpenStudy (anonymous):

Sorry I missed the square root \[ \sqrt{6x-8}=x\] squaring both sides \[ 6x-8=x^2\] \[ x^2 -6x +8=0\] Now solve

OpenStudy (anonymous):

Thank you ! @sangya21

OpenStudy (anonymous):

Anytime. )